# A compact QUBO encoding of computational logic formulae demonstrated on cryptography constructions

### v1

@@ 4.2 Encoding constructions of MD5, SHA-1 and SHA-256 hash functions
We also provide QUBO instances on larger hash function like the SHA1 and SHA-256. The main difference compared to MD5 in relation to QUBO encoding are the implemented modular additions. While MD5 incorporates one addition over $4$ inputs and one addition on $2$ inputs, SHA involves additions that can be joined into a single addition with $5$ inputs, while the additions in SHA-256 can be grouped into additions with $6$ and $7$ inputs. To encode these addition into a QUBO form, we further modify Eq. (48). Aside having more inputs, we also need to introduce one more carry bit in each block accounting for possible overflows between the blocks. Because of the extra carry bit, the magnitude of the maximal elements in the QUBO matrix gets increased to $\sim 2^{2B+4}$ as reported in Table. 3 for the encoding of SHA1a and SHA-256a for $B=6$. Comparing to the QUBO encoding reported in Ref. [36] on SHA-256, we provide a QUBO encoding with $30255$ variables compared to $47808$ variables of Ref. [36]. This was mainly achieved by partitioning the modular additions into blocks of size $B=6$. When targeting QUBO instances with smaller elements, we can use blocks with $B=1$, still improving on the result of Ref. [36] with $42899$ QUBO variables and maximal elements of $55$. Our results can be accesses via GitHub repository at Ref. [75].
@@ 5 Conclusions
The most important limitations or unknown aspects of QUBO encoding is when dealing with general binary functions constructed for special functions. In circuit minimization and optimization, though finding satisfiable possibilities scales even higher, unsatisfiability typically becomes intractable at $6$-variable functions, due to the practical explosion in the search space [70]. While techniques like symmetry breaking, slack variables, guided counter-example optimization, and other common reductions help, these improvements generally have only served to make e.g. $6$-variable functions tractable for many contexts. With ILPs, the situation is very similar as expected. As finding the optimal encoding for larger boolean functions (like the sum-of-products formulae and integer multiplication) would significantly improve on the variable count in QUBO encodings, the study on further improvement of the general ILP formulation of binary functions and solving these problems deserves significant attention. We leave these studies for a future work. Future patent applications will be filed to further develop and expand the encoding techniques described in this paper.

### v2

@@ 4.2 Encoding constructions of MD5, SHA-1 and SHA-256 hash functions
We also provide QUBO instances on larger hash function like the SHA1 and SHA-256. The main difference compared to MD5 in relation to QUBO encoding are the implemented modular additions. While MD5 incorporates one addition over 44 inputs and one addition on 22 inputs, SHA involves additions that can be joined into a single addition with 55 inputs, while the additions in SHA-256 can be grouped into additions with 66 and 77 inputs. To encode these addition into a QUBO form, we further modify Eq. (48). Aside having more inputs, we also need to introduce one more carry bit in each block accounting for possible overflows between the blocks. Because of the extra carry bit, the magnitude of the maximal elements in the QUBO matrix gets increased to $\sim 2^{2B+4}$ as reported in Table. 3 for the encoding of SHA1a and SHA-256a for $B=6$. Comparing to the QUBO encoding reported in Ref. [36] on SHA-256, we provide a QUBO encoding with $30255$ variables compared to $47808$ variables of Ref. [36]. This was mainly achieved by partitioning the modular additions into blocks of size $B=6$. When targeting QUBO instances with smaller elements, we can use blocks with $B=1$, still improving on the result of Ref. [36] with $42899$ QUBO variables and maximal elements of $55$. Our results can be accesses via GitHub repository at Ref. [75].
@@ 5 Conclusions
The most important limitations or unknown aspects of QUBO encoding is when dealing with general binary functions constructed for special functions. In circuit minimization and optimization, though finding satisfiable possibilities scales even higher, unsatisfiability typically becomes intractable at 66-variable functions, due to the practical explosion in the search space [70]. While techniques like symmetry breaking, slack variables, guided counter-example optimization, and other common reductions help, these improvements generally have only served to make e.g. 66-variable functions tractable for many contexts. With ILPs, the situation is very similar as expected. As finding the optimal encoding for larger boolean functions (like the sum-of-products formulae and integer multiplication) would significantly improve on the variable count in QUBO encodings, the study on further improvement of the general ILP formulation of binary functions and solving these problems deserves significant attention. We leave these studies for a future work. Future patent applications will be filed to further develop and expand the encoding techniques described in this paper.